Zeronknight (zeronknight):

Choose a solution to the following inequality. *only one value is correct 2x + 9 < 15 A. 5 B. 7 C. 3 D. 1

OpenStudy (kendricklamar2014):

\[2x + 9 < 15\] Subtract 9 from both sides: \[2x+9−9<15−9\] \[= 2x<6\] Divide both sides by 2: \[\frac{ 2x }{ 2 } < \frac{ 6 }{ 2 }\] \[x < 3\]

Zeronknight (zeronknight):

So 2 would be negative?

OpenStudy (kendricklamar2014):

There are no negative numbers..

Zeronknight (zeronknight):

Forgot you are adding 9 to 2 mb.

OpenStudy (kendricklamar2014):

Its ok, im here to help you understand :)